# Split Gibbs Discrete Diffusion Posterior Sampling

### v2

@@ A.1 Lemmas
Consider the function $g(z)=1-z-e^{-z}+(u-1)(1-z)$. When $u\geq 0$, this function is maximized when $z=-\log(1+u)$. By Assumption (i) and using $u=\tilde{\bm{Q}}_{t}^{[j,i]}/{\bm{Q}}_{t}^{[j,i]}-1$,
$\sup_{z_{ij}}\Big{(}\frac{\tilde{\bm{Q}}_{t}^{[j,i]}}{{\bm{Q}}_{t}^{[j,i]}}(1-z_{ij})-e^{-z_{ij}}\Big{)}=(1+u)\ln(u+1)\leq|u|\leq\epsilon,\ \text{for}\ |u|\leq 1$ (31)
Combining with Eq. 30, we have
@@ A.2 Proof of Theorem 1
Consider discretizing a prior step $[0,t^{*}]$ uniformly into $H$ intervals, $[h\delta,(h+1)\delta]$, for $h=0,\dots,H-1$, where $\delta=t^{*}/H$. In the $h$-th interval,
@@
$\partial_{t}\mathsf{KL}(\mu_{t}\|\pi_{t})\leq-\frac{1}{2}\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})+\frac{2}{c}\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\tilde{\bm{Q}}_{h\delta}\mu_{t})$ (34)
@@
Finally, taking summation over $k=0,\dots,K-1$, and applying Lemma 1 for each likelihood step,
@@ A.3 Potential function of split Gibbs samplers
As defined in Eq. 7, split Gibbs samplers draws samples from the augmented distribution
@@
Also, the similar derivation holds for $\pi^{Z}({\mathbf{z}};\eta)$. This result has also been shown in [43]. Combining this with Theorem 1, SGDD is guaranteed to sample from the true posterior distribution when $\eta$ goes to zero.
@@ Appendix B Ablation studies
As shown in Fig. 5, running SGDD with the annealing noise schedule achieves the highest PSNR and also the fastest convergence speed. Moreover, when $\eta$ is large, SGDD converges quickly to a poor solution, whereas when $\eta$ is small, it converges slowly to an accurate solution. This finding aligns with our intuition that SGDD is most accurate with $\eta\to 0$, while sampling with $\eta>0$ regularizes and simplifies the sampling problem.
@@
Theoretical insights. Since Eq. 41 holds when $\eta\to 0$, a natural question is why we cannot fix a small $\eta$ to begin with.
The answer to this question is contained in Theorem 1. When fixing a small $\eta$, we also fix a small $t^{*}$ as $\sigma(t^{*})=\eta$. Then the transition matrix ${\bm{Q}}_{\tau}$ ($\tau\in[0,t^{*}]$) will be ill-posed, in the sense that ${\bm{Q}}_{\tau}^{[i,j]}$ is either very large or very close to 0. In this case, the relative Fisher information $\mathsf{FI}_{{\bm{Q}}_{\tau}}(\mu_{\tau}\|\pi_{\tau})$ becomes a poor criterion of measuring the divergence between $\mu_{\tau}$ and $\pi_{\tau}$. To see this, we note that the spectral gap $\gamma({\bm{Q}}_{\tau})$ has
@@
We run SGDD with different configurations to evaluate its performance under different computing budgets. We control the number of iterations, $K$, and the number of unconditional sampling steps $T$ in each prior sampling step, so the number of function evaluations (NFE) of the discrete diffusion model is $KT$. We evaluate SGDD with NFEs ranging from $100$ to $2$k. The detailed configuration is specified in Table 6.
We present a plot in Fig. 5 showing the tradeoff between sample quality and computing budgets in terms of NFEs. SGDD is able to generate high-quality samples with a reasonable budget comparable to unconditional generation. Note that we calculate the NFEs for SVDD and SMC with the number of Monte Carlo samples $\mathrm{mc}=1$ and the number of Euler steps $H=1$, while some results in our experiments are done with larger $\mathrm{mc}$ and $H$ for better performance. We report both the number of total NFEs and the number of sequential NFEs that decide the cost of the algorithm when it is fully parallelized.
@@
Configuration ##\## Iterations Prior sampling steps NFE (total/sequential) Runtime (sec/sample)
DPS $100$ $102400/100$ 820
SVDD-PM ($M=20$) $128$ $2560/128$ 16
SMC $(M=20)$ $128$ $5120/256$ 21
SGDD-$50$ $25$ $2$ $50$ 4
SGDD-$100$ $25$ $4$ $100$ 6
SGDD-$200$ $40$ $5$ $200$ 8
@@
We evaluate the value of discrete diffusion priors in solving discrete inverse problems by comparing our method to the Metropolis-Hastings algorithm without a diffusion prior. We pick the AND problem on discretized MNIST as an example, in which we choose $\gamma D$ random pairs of pixels and compute their AND values as measurement ${\mathbf{y}}$. Fig. 6 shows the reconstructed samples of SGDD and direct MCMC sampling under different levels of measurement sparsity. As demonstrated in the figure, the discrete diffusion model complements the missing information when the measurement is sparse (e.g., $\gamma=2,4,8$), underscoring the importance of discrete diffusion priors in SGDD for solving discrete inverse problems.
@@ C.2.1 DPS
However, DPS is not directly applicable to inverse problems in discrete-state spaces since propagating gradients through ${\bm{G}}$ and ${\bm{D}}_{\theta}$ in Eq. 47 is impossible. Therefore, we consider the counterpart of DPS in discrete spaces. We modify the continuous-time Markov chain of the discrete diffusion model by
@@ C.3 Hyperparameters
We use an annealing noise schedule of $\eta_{k}=\eta_{\min}^{k/K}\eta_{\max}^{1-k/K}$ with $\eta_{\min}=10^{-4}$ and $\eta_{\max}=20$. We run SGDD for $K$ iterations. In each likelihood sampling step, we run Metropolis-Hastings for $T$ steps, while in each prior sampling step, we run a few-step Euler discrete diffusion sampler with $H$ steps. The hyperparameters used for each experiment are listed in Table 7. We also include the dimension of data spaces $\mathcal{X}^{D}$ for each experiment in Table 7, where $|\mathcal{X}|=N$.
@@ Appendix D Additional Experimental Results
Specifically, we define the posterior distribution as proportional to $p({\mathbf{x}})p(-\beta\ell({\bm{G}}({\mathbf{x}}),{\mathbf{y}}))$, where $\ell$ is the cross-entropy loss between the predicted class probability and target class label ${\mathbf{y}}$. We let ${\mathbf{y}}=0$ to guide diffusion models to generate digits 00, and extract all digits 00 in the dataset as the approximated true posterior distribution. We train another classifier with a different random seed to evaluate the generated samples. To measure the divergence of the generated samples from the posterior distribution, we extract the features of MNIST before the last layer of a CNN classification network, and compute the multi-scale intrinsic distance (MSID) score.
Experimental results in Table 9 show that SGDD effectively samples from the true posterior distribution, as it achieves a low MSID score and a high class accuracy of $98.5\%$.

### v3

@@ A.1 Lemmas
Consider the function $g(z)=-uz-e^{-z}$. When $u\geq 0$, this function is maximized when $z=-\log u$. Using $u={\bm{Q}}_{t}^{[j,i]}/\tilde{\bm{Q}}_{t}^{[j,i]}$,
$\sup_{z_{ij}}\Big(-\frac{{\bm{Q}}_{t}^{[j,i]}}{\tilde{\bm{Q}}_{t}^{[j,i]}}z_{ij}-e^{-z_{ij}}\Big)=(u-1)\log u\leq|u-1|\leq\epsilon,$ (31)
for $\epsilon\leq 1$. Combining with Eq.˜30, we have
@@
Additional Notes on Lemma 3.
With a slightly different derivation, assumptions (i) and (v) can potentially be replaced by the following assumption that links directly to the score entropy loss defined in Definition 3.1 of Lou et al. [29].
Alternative assumption.
  • (i)
    Score function is well estimated, i.e., the score entropy loss for distributions on the stationary process is bounded:
    $\mathbb{E}_{{\mathbf{x}}_{i}\sim\pi_{t}}\mathsf{SE}_{{\bm{Q}}^{\mathsf{fw}}}({\mathbf{x}}_{i};\theta,t):=\mathbb{E}_{{\mathbf{x}}_{i}\sim\pi_{t}}\sum_{{\mathbf{x}}_{j}\neq{\mathbf{x}}_{i}}K\left(\frac{s_{\theta}({\mathbf{x}}_{i};t)_{{\mathbf{x}}_{j}}}{s({\mathbf{x}}_{i};t)_{{\mathbf{x}}_{j}}}\right){\bm{Q}}_{t}^{\mathsf{fw}\ [i,j]}s({\mathbf{x}}_{i};t)_{{\mathbf{x}}_{j}}\leq\epsilon_{\mathsf{SE}},$
    for any $t\in[0,1]$, where $K(a)=a-1-\log a$.
Lemma 4.
When the learned score function $s_{\theta}$ satisfies Assumption (i),
$\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},{\bm{Q}}_{t}\pi_{t})\leq\epsilon_{\mathsf{SE}}.$ (33)
Proof.
$\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},{\bm{Q}}_{t}\pi_{t})$ $=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\pi_{t}({\mathbf{x}}_{i})\tilde{\bm{Q}}_{t}^{[j,i]}\left(\frac{{\bm{Q}}_{t}^{[j,i]}}{\tilde{\bm{Q}}_{t}^{[j,i]}}(1-z_{ij})-e^{-z_{ij}}\right)\right]+\sum_{{\mathbf{x}}_{i}}\pi_{t}({\mathbf{x}}_{i})({\bm{Q}}_{t}^{[i,i]}-\tilde{\bm{Q}}_{t}^{[i,i]})$
$\leq\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\pi_{t}({\mathbf{x}}_{i})\tilde{\bm{Q}}_{t}^{[j,i]}\frac{s({\mathbf{x}}_{i};t)}{s_{\theta}({\mathbf{x}}_{i};t)}\log\frac{s({\mathbf{x}}_{i};t)}{s_{\theta}({\mathbf{x}}_{i};t)}\right]-\sum_{{\mathbf{x}}_{i}}\pi_{t}({\mathbf{x}}_{i})\sum_{j\neq i}({\bm{Q}}_{t}^{[j,i]}-\tilde{\bm{Q}}_{t}^{[j,i]})$ (34)
$=\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\pi_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}\log\frac{s({\mathbf{x}}_{i};t)}{s_{\theta}({\mathbf{x}}_{i};t)}\right]-\sum_{{\mathbf{x}}_{i}}\pi_{t}({\mathbf{x}}_{i})\sum_{j\neq i}({\bm{Q}}_{t}^{[j,i]}-\tilde{\bm{Q}}_{t}^{[j,i]})$
$=\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\pi_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{\mathsf{fw}[i,j]}\Big(s_{\theta}({\mathbf{x}}_{i};t)_{{\mathbf{x}}_{j}}\log\frac{s({\mathbf{x}}_{i};t)}{s_{\theta}({\mathbf{x}}_{i};t)}-s({\mathbf{x}}_{i};t)+s_{\theta}({\mathbf{x}}_{i};t)\Big)\right]$
$=\mathbb{E}_{{\mathbf{x}}_{i}\sim\pi_{t}}\sum_{{\mathbf{x}}_{j}\neq{\mathbf{x}}_{i}}\left[K\Big(\frac{s_{\theta}({\mathbf{x}}_{i};t)}{s({\mathbf{x}}_{i};t)}\Big){\bm{Q}}_{t}^{\mathsf{fw}[i,j]}s({\mathbf{x}}_{i};t)_{{\mathbf{x}}_{j}}\right]\leq\epsilon_{\mathsf{SE}}.$ (35)
where Eq.˜34 comes from the previous Lemma 3. ∎
@@ A.2 Proof of Theorem 1
Consider discretizing a prior step $0,t^{*}]$ uniformly into $H$ intervals, $h\delta,(h+1)\delta]$, for $h=0,\dots,H-1$, where $\delta=t^{*}/H$. In the $h$-th interval,
@@
$\partial_{t}\mathsf{KL}(\pi_{t}\|\mu_{t})\leq-\frac{1}{2}\mathsf{FI}_{\tilde{\bm{Q}}_{h\delta}}(\pi_{t}\|\mu_{t})+\frac{2}{c}\mathbb{E}_{{\mathbf{x}}_{i}\sim\pi_{t}}\mathcal{L}_{\tilde{\bm{Q}}_{h\delta}}(\pi_{t},{\bm{Q}}_{t}\pi_{t}).$ (37)
@@
Finally, taking summation over $k=0,\dots,K-1$, and applying Lemma˜1 for each likelihood step,
@@ A.3 Potential function of split Gibbs samplers
As defined in Eq.˜7, split Gibbs samplers draws samples from the augmented distribution
@@
Also, the similar derivation holds for $\pi^{Z}({\mathbf{z}};\eta)$. This result has also been shown in [46]. Combining this with Theorem˜1, SGDD is guaranteed to sample from the true posterior distribution when $\eta$ goes to zero.
@@ Appendix B Ablation studies
As shown in Fig.˜5, running SGDD with the annealing noise schedule achieves the highest PSNR and also the fastest convergence speed. Moreover, when $\eta$ is large, SGDD converges quickly to a poor solution, whereas when $\eta$ is small, it converges slowly to an accurate solution. This finding aligns with our intuition that SGDD is most accurate with $\eta\to 0$, while sampling with $\eta>0$ regularizes and simplifies the sampling problem.
@@
Theoretical insights. Since Eq.˜44 holds when $\eta\to 0$, a natural question is why we cannot fix a small $\eta$ to begin with.
The answer to this question is contained in Theorem˜1. When fixing a small $\eta$, we also fix a small $t^{*}$ as $\sigma(t^{*})=\eta$. Then the transition matrix ${\bm{Q}}_{\tau}$ ($\tau\in[0,t^{*}]$) will be ill-posed, in the sense that ${\bm{Q}}_{\tau}^{[i,j]}$ is either very large or very close to 0. In this case, the relative Fisher information $\mathsf{FI}_{{\bm{Q}}_{\tau}}(\mu_{\tau}\|\pi_{\tau})$ becomes a poor criterion of measuring the divergence between $\mu_{\tau}$ and $\pi_{\tau}$. To see this, we note that the spectral gap $\gamma({\bm{Q}}_{\tau})$ has
@@
We run SGDD with different configurations to evaluate its performance under different computing budgets. We control the number of iterations, $K$, and the number of unconditional sampling steps $T$ in each prior sampling step, so the number of function evaluations (NFE) of the discrete diffusion model is $KT$. We evaluate SGDD with NFEs ranging from $100$ to 22k. The detailed configuration is specified in Table˜6.
We present a plot in Fig.˜5 showing the tradeoff between sample quality and computing budgets in terms of NFEs. SGDD is able to generate high-quality samples with a reasonable budget comparable to unconditional generation. Note that we calculate the NFEs for SVDD and SMC with the number of Monte Carlo samples $\mathrm{mc}=1$ and the number of Euler steps $H=1$, while some results in our experiments are done with larger $\mathrm{mc}$ and $H$ for better performance. We report both the number of total NFEs and the number of sequential NFEs that decide the cost of the algorithm when it is fully parallelized.
@@
Configuration #\# Iterations Prior sampling steps NFE (total/sequential) Runtime (sec/sample)
DPS $100$ - $102400/100$ 820
SVDD-PM ($M=20$) $128$ - $2560/128$ 16
SMC $(M=20)$ $128$ - $5120/256$ 21
SGDD-$50$ $25$ 22 $50$ 4
SGDD-$100$ $25$ 44 $100$ 6
SGDD-$200$ $40$ 55 $200$ 8
@@
We evaluate the value of discrete diffusion priors in solving discrete inverse problems by comparing our method to the Metropolis-Hastings algorithm without a diffusion prior. We pick the AND problem on discretized MNIST as an example, in which we choose $\gamma D$ random pairs of pixels and compute their AND values as measurement $\mathbf{y}$}. Fig.˜6 shows the reconstructed samples of SGDD and direct MCMC sampling under different levels of measurement sparsity. As demonstrated in the figure, the discrete diffusion model complements the missing information when the measurement is sparse (e.g., $\gamma=2,4,8$), underscoring the importance of discrete diffusion priors in SGDD for solving discrete inverse problems.
@@ C.2.1 DPS
However, DPS is not directly applicable to inverse problems in discrete-state spaces since propagating gradients through $\bm{G}$} and ${\bm{D}}_{\theta}$ in Eq.˜50 is impossible. Therefore, we consider the counterpart of DPS in discrete spaces. We modify the continuous-time Markov chain of the discrete diffusion model by
@@ C.3 Hyperparameters
We use an annealing noise schedule of $\eta_{k}=\eta_{\min}^{k/K}\eta_{\max}^{1-k/K}$ with $\eta_{\min}=10^{-4}$ and $\eta_{\max}=20$. We run SGDD for $K$ iterations. In each likelihood sampling step, we run Metropolis-Hastings for $T$ steps, while in each prior sampling step, we run a few-step Euler discrete diffusion sampler with $H$ steps. The hyperparameters used for each experiment are listed in Table˜7. We also include the dimension of data spaces $\mathcal{X}^{D}$ for each experiment in Table˜7, where $|\mathcal{X}|=N$.
@@ Appendix D Additional Experimental Results
Specifically, we define the posterior distribution as proportional to $p({\mathbf{x}})p(-\beta\ell({\bm{G}}({\mathbf{x}}),{\mathbf{y}}))$, where $\ell$ is the cross-entropy loss between the predicted class probability and target class label $\mathbf{y}$}. We let ${\mathbf{y}}=0$ to guide diffusion models to generate digits 0, and extract all digits 0 in the dataset as the approximated true posterior distribution. We train another classifier with a different random seed to evaluate the generated samples. To measure the divergence of the generated samples from the posterior distribution, we extract the features of MNIST before the last layer of a CNN classification network, and compute the multi-scale intrinsic distance (MSID) score.
Experimental results in Table˜9 show that SGDD effectively samples from the true posterior distribution, as it achieves a low MSID score and a high class accuracy of $98.5\%$.